OpenStudy (anonymous):

HELP 2 Math questions timed quiz!!!! Distribute: 4(3x + 7) (Points : 1) 7x + 7 12x + 7 7x + 28 12x + 28 OTHERS I AM TYPING BELOW

OpenStudy (anonymous):

\[Distribute \frac{ 1 }{ 2 } (8x16)\]

OpenStudy (anonymous):

16x- 32 4x-8 16x-8 4x-16

OpenStudy (anonymous):

4x-16 = x=4

OpenStudy (anonymous):

16x-32=2

OpenStudy (anonymous):

4x-8 x=2

OpenStudy (anonymous):

16x-8 x=2

OpenStudy (anonymous):

no thoughs are the answers

OpenStudy (anonymous):

those ***

OpenStudy (anonymous):

try re-submitting this in Algebra and quickly!!!!

OpenStudy (anonymous):

click find more subjects @ the top of page

OpenStudy (anonymous):

2 over from the message icon

OpenStudy (anonymous):

oh this was math at one point oops!!! this thing has been messing up all day

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

hurry:)